# Some Remarks on Positive/Negative Feedback

### v1

@@ 4 Conclusion
The claim in this note is purely existential in nature in the sense that a commonly-perceived dichotomy of negative and positive feedback is shown not to be valid in general. This has been done by proving the existence of a nonlinear counter-example: the practical utility and quantitative behaviour of the feedback employed therein has not been addressed. Whilst the counter-example is open to a criticism of artificiality, it does, however, serve to illustrate that the dichotomy “negative/positive feedback” equates to “good/bad behaviour” may be inaccurate. The counter-example is based on a wider body of work in the area of “funnel control” (a terminology consistent with Figure 1). The interested reader is referred to [2] for a review of this area.

### v2

@@ 5 Conclusion
A dichotomy of negative and positive feedback has been shown to be invalid in general. This has been done by establishing the existence of a nonlinear example which serves to illustrate that “negative/positive feedback” need not equate to “good/bad behaviour”. In this sense, the claim in this note is purely existential in nature. In essence, it simply provides an example in which a feedback and its reversed-polarity counterpart both cause “good” behaviour. The practical utility and quantitative behaviour of the feedback employed therein have not been addressed. The particular example in Section 3 may be open to a criticism of artificiality: we stress that its role is only to show that a system of the general form (1.1) can have non-unique control direction, thereby ensuring the existence of a stabilizing feedback with the property that its reversed-polarity counterpart is also stabilizing.